OpenStudy (anonymous):

Solve the system. y = 4x + 12 and y = -2x + 6

OpenStudy (turingtest):

pick your method, substitution or elimination?

OpenStudy (anonymous):

substitution looks ok since it only 2 equations and 2 variables and you are already solved for y.

OpenStudy (anonymous):

4x + 12 = -2x + 6

OpenStudy (anonymous):

-1,3 ?

mathslover (mathslover):

Did you mean to say that, x = -1 and y = 3 @IllGiveYouMyAll ?

OpenStudy (anonymous):

Yes, is that correct? @mathslover

mathslover (mathslover):

x = -1 seems okay. But, y = 3 does not ...

mathslover (mathslover):

How did you get to the above conclusion ?

OpenStudy (anonymous):

I guessed. It's so confusing... Ugh.

mathslover (mathslover):

See, is it clear to you , that : 4x + 12 = -2x + 6 ?

OpenStudy (anonymous):

Well yes, but once i start solving it i seem to mess up.

mathslover (mathslover):

No problem. We're here to help you with that. Okay. Let's start solving it. First of all, \(\boxed{4x + 12 = -2x + 6}\) Now, add \(\bf{2x}\) both sides. \(\implies 4x + 12 + \underline{2x} = -2x + \underline {2x} + 6\) [Notice the underlined "2x" there]

mathslover (mathslover):

Now, can you tell me, what is : 4x + 12 + 2x ?

OpenStudy (anonymous):

6x + 12?

mathslover (mathslover):

Good :) . Now, similarly solve for : -2x + 2x + 6

OpenStudy (anonymous):

6?

mathslover (mathslover):

Very good ! So, you have now : 6x + 12 = -6 Can you solve for "x" from the above equation now?

OpenStudy (anonymous):

3?

mathslover (mathslover):

No , no! A little bit mistake. 6x + 12 = -6 Add 6 both sides 6x + 12 + 6 = -6 + 6 6x + 18 = 0 Subtract 18 both sides 6x + 18 - 18 = 0 - 18 6x = -18 Divide 6 both sides \(\cfrac{6x}{6} = \cfrac{-18}{6} \) x = -3

OpenStudy (anonymous):

Oh, oh i see. I never know if when you multiply if it will be negative or positive.

mathslover (mathslover):

Okay, now you have x = -3. You had in first, y = 4x + 12

mathslover (mathslover):

Put the value of x = -3 in : y = 4x + 12 and solve for y

OpenStudy (anonymous):

0?

mathslover (mathslover):

Yes, but wait... a mistake there.

mathslover (mathslover):

I did a mistake there . Sorry, : 6x + 12 = 6 [Not -6]

mathslover (mathslover):

i.e. : \(6x + 12 = 6\) Sorry for the blunder.

mathslover (mathslover):

Though, process will be exactly same. 6x + 12 = 6 Subtract 12 both sides 6x + 12 -12 = 6 - 12 6x = -6 x = -1

mathslover (mathslover):

y = 4x + 12 [Given] y = 4(-1) + 12 Now solve for y.

OpenStudy (anonymous):

So the answer is (-1,-3)?

mathslover (mathslover):

No no! y = 4(-1) + 12 y = -4 + 12 y = ?

OpenStudy (anonymous):

8!

mathslover (mathslover):

Yep, good. So, y = 8 and the solution is : (-1,8)

OpenStudy (anonymous):

Thank you SO much! c:

mathslover (mathslover):

You're very welcome :)

mathslover (mathslover):

And also, \(\Huge\bf \color{black}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{black}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{#ff7100}{Welcome~to~\color{#00B4ff}{\underline{\overline{Open}}}\color{#7cc517}{\underline{\overline{Study!!}}}}\)